OpenStudy (anonymous):

how do you convert the equation of a line in 3 dimensions to parametric

OpenStudy (anonymous):

What types of equation are you starting with? Is it a vector equation or a Cartesian equation?

OpenStudy (anonymous):

vector equations

OpenStudy (anonymous):

So you're starting from r = (ai+bj+ck) + p(mi+nj+qk). This can be written as r=(a+pm)i + (b+pn)j + (c+pq)k Hence the parametric equation would be x=a+pm, y=b+pn, z=c+pq

OpenStudy (anonymous):

where p would be the parameter

OpenStudy (anonymous):

p is like the t?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have to factor it out of the given equation then?

OpenStudy (anonymous):

in the vector equation you usually have lambda, which I called p here. That p would become the t for the parametric form.